# https://semantic-gaussians.github.io

### v1

@@ Abstract
Open-vocabulary 3D scene understanding presents a significant challenge in computer vision, with wide-ranging applications in embodied agents and augmented reality systems. Previous approaches have adopted Neural Radiance Fields (NeRFs) to analyze 3D scenes. In this paper, we introduce Semantic Gaussians, a novel open-vocabulary scene understanding approach based on 3D Gaussian Splatting. Our key idea is distilling pre-trained 2D semantics into 3D Gaussians. We design a versatile projection approach that maps various 2D semantic features from pre-trained image encoders into a novel semantic component of 3D Gaussians, without the additional training required by NeRFs. We further build a 3D semantic network that directly predicts the semantic component from raw 3D Gaussians for fast inference. We explore several applications of Semantic Gaussians: semantic segmentation on ScanNet-20, where our approach attains a 4.2% mIoU and 4.0% mAcc improvement over prior open-vocabulary scene understanding counterparts; object part segmentation, scene editing, and spatial-temporal segmentation with better qualitative results over 2D and 3D baselines, highlighting its versatility and effectiveness on supporting diverse downstream tasks.
Keywords:
Open-vocabulary Scene Understanding 3D Gaussian Splatting
footnotetext: * Equal contribution.footnotetext: † Corresponding authors.
@@ 1 Introduction
Various methods have been proposed to achieve open-vocabulary 3D scene understanding, relying on different 3D scene representations such as multi-view RGB images [11], point clouds [27, 33], and Neural Radiance Fields (NeRFs) [21, 15]. Approaches based on these representations have their pros and cons: Multi-view images are the most straightforward representation of 3D scenes, but allowing open-vocabulary understanding usually involves 2D vision-language models [1], which could struggle with consistency across different views, likely due to a lack of visual geometric knowledge; Point clouds are popular and well-studied, but the inherent sparsity nature of point clouds limits the application of open-vocabulary scene understanding upon them  [27], e.g., it is challenging to obtain a dense prediction on a 2D view; injecting open-vocabulary semantics to NeRFs could enjoy both high 2D rendering quality and dense free-form visual recognition [15], but the implicit design requires open-vocabulary recognition training for every new scene. A recent alternative scene representation is 3D Gaussian Splatting proposed by Kerbl et al. [14]. It utilizes 3D Gaussian points with color, opacity, and covariance matrix to represent the 3D scene, which can be learned from multi-view RGB images via gradient descent training. It attains the NeRF-level view rendering quality while preserving the explicit point-based characteristics similar to point clouds, making it suitable for open-vocabulary 3D scene understanding.
@@
In this work, we propose Semantic Gaussians, a novel approach to open-vocabulary 3D scene understanding building upon the benefits of 3D Gaussian Splatting. The core idea of Semantic Gaussians is to distill the knowledge from pre-trained 2D encoders into 3D Gaussians, thereby assigning a semantic component to each Gaussian point. To achieve this, we establish correspondence between 2D pixels and 3D Gaussian points and propose a versatile projection framework to map the semantic features of 2D pixels onto each 3D Gaussian point. Our framework is rather flexible and can leverage arbitrary pre-trained 2D models, such as OpenSeg [10], CLIP [29], VLPart [32], etc., to generate pixel-wise semantic features on 2D RGB images. Compared to the NeRF-based approach [15, 36], our method injects semantic components into 3D Gaussians without additional training, allowing for effective open-vocabulary scene queries.
@@
In addition to projection, we further introduce a 3D semantic network that directly predicts open-vocabulary semantic components out of raw 3D Gaussians. Specifically, we employ MinkowskiNet [5], a 3D sparse convolution network to process 3D Gaussians. The 3D convolution network takes raw RGB Gaussians as input and is supervised by the semantic components of Gaussians obtained from the aforementioned projection method. As a result, we may simply run this network to obtain the semantic components, enabling faster inference. Note that the prediction of the 3D semantic network can be combined with the projected features to further improve the quality of semantic components in Gaussians and open-vocabulary scene understanding performances.
@@
We conduct experiments on the ScanNet semantic segmentation benchmark [6], and prove our efficiency compared to 2D pre-trained models. Besides semantic segmentation, we also explore diverse applications of Semantic Gaussians, including 3D part segmentation on the MVImgNet object dataset [35], scene editing in multi-object scenes, and spatiotemporal tracking on 4D dynamic Gaussians [24].
@@
    We conduct experiments on the ScanNet segmentation benchmark to demonstrate the effectiveness of our method on open-vocabulary scene understanding and explore various applications including object part segmentation, scene editing, and spatiotemporal tracking.
2 Related Work
2.1 3D Scene Representation
@@ 2 Related Work
Recently, Kerbl et al. [14] have proposed a new novel view synthesis method called 3D Gaussian Splatting, which represents the 3D scene with a set of 3D Gaussians. This method has demonstrated real-time rendering capabilities at 1080p resolution, achieving a remarkable 60 frames per second while maintaining state-of-the-art visual quality. The innovation behind 3D Gaussian Splatting lies in its incorporation of point-based $\alpha$-blending and a differentiable tile rasterizer, enabling efficient rendering. Though it obtains a dense set of Gaussians via optimization, all parameters in these 3D Gaussians are explicit and editable. The speed enhancement and explicit parameterization position 3D Gaussian Splatting as a highly promising representation method. Building upon its success in novel view synthesis, Luiten et al. [24] extended the concept to Dynamic 3D Gaussians, explicitly modeling 3D Gaussians at different time steps to accommodate 4D dynamic scenes. Furthermore, Chen et al. [4] proposed a text-to-3D generation technique leveraging 3D Gaussian Splatting. In this study, we propose an open-vocabulary 3D scene understanding method, leveraging the advantages offered by 3D Gaussian Splatting.
2.2 Open-Vocabulary Scene Understanding
@@
Scene understanding from 2D.   Encouraged by the availability of adequate text-image datasets and the advancement in vision language models, the field of 2D open-vocabulary scene understanding has made significant progress in recent years. Prevailing approaches [20, 25, 18] distill knowledge from large-scale pre-trained foundation models (e.g., CLIP [29]) to achieve zero-shot understanding, including recognizing long-tail objects and understanding synonymous labels. However, these methods are limited to small partial scenes represented by a single 2D image. When it comes to 3D scenes, the prediction result of these 2D models can hardly remain consistent between different angles of view. In contrast, our work relies on these 2D pre-trained models to achieve 3D scene understanding, segmenting, and understanding the scene from a panoptic perspective. Moreover, the proposed 3D network can perform 3D-only scene understanding in the absence of 2D pre-trained models and images.
@@
Scene understanding from 3D.  Open-vocabulary 3D scene understanding has been a long-standing challenge in computer vision. Some point-cloud-based methods[27, 12, 7] encode the semantic features from 2D pre-trained models into 3D scene points to achieve open-vocabulary 3D scene understanding. To achieve high-quality rendering and scene understanding simultaneously, feature field distillation in NeRF has been well explored. Early works such as Semantic-NeRF [36] and Panoptic Lifting [31] embed semantic labels into NeRF, resulting in precise 3D segmentation maps. Encouraged by this idea, another branch of methods such as Distill Feature Fields [17], NeRF-SOS [8], LERF [15], 3D-OVS [21] integrate pixel-wise semantic embeddings from pre-trained models such as LSeg [18], CLIP or DINO [3] into NeRFs, achieving open-vocabulary 3D scene understanding. Our work shares a similar idea with those NeRF-based methods, while the Semantic Gaussians requires no extra training for Gaussians. The explicit nature of 3D Gaussian Splatting enables Semantic Gaussians to achieve versatile projection from 2D semantic maps into 3D Gaussian points.
3 Semantic Gaussians
@@
In this section, we illustrate the framework of our Semantic Gaussians. Fig. 2 depicts the overall framework of our method. Semantic Gaussians starts from a group of 3D Gaussians, performing scene understanding on it through 2D versatile projection and 3D semantic network processing. We first introduce our versatile projection method that projects 2D semantic embeddings from various pre-trained vision-language models into 3D Gaussian points (Sec. 3.1). We then depict the 3D semantic network that learns from the projected features and predicts the semantics of 3D Gaussians in unseen scenes (Sec. 3.2). At last, we describe the feature ensemble process and use the prediction result to support various applications (Sec. 3.3).
3.1 2D Versatile Projection
@@
3.1.2 Unifiying Various 2D Features with SAM.
Accommodating a variety of 2D pre-trained features is non-trivial as they can be pixel-level segmentation network (e.g., OpenSeg [10], LSeg [18]), instance-level recognition network (e.g., GroundingDINO [22], VLPart [32]), or image-level classification network (e.g., CLIP [29]). Semantic Gaussians utilizes Segment Anything (SAM) [16] to produce fine segmentation maps for each model. For pixel-level models, SAM can refine the segmentation boundary. Given an RGB image $\mathbf{I}$, we use everything prompt in SAM to generate $N$ binary masks $\mathbf{M}_{1},\cdots,\mathbf{M}_{N}$. We calculate the average pooling of embeddings in each mask $\mathbf{M}_{i}$, and assign it as the embedding of all pixels in this mask: $s[\mathbf{M}_{i}]=AvgPool(s[\mathbf{M}_{i}])$. For instance-level models, we use SAM as the postprocessing module to get fine masks. Similar to Grounded-SAM [30], we use the prediction result from the pre-trained model as the box prompt of SAM. After getting the binary mask, we assign the CLIP embedding of this instance to all the pixels within the instance region. For image-level models, SAM can be a preprocessing module to get region proposals. We use the “everything” prompt in SAM to get various proposed regions. Each region is padded, cropped, and resized to $224\times 224$, and fed into the image-level model to get semantic embeddings. Similarly, the semantic embeddings are assigned to all pixels within each proposed region.
3.1.3 2D-3D Projection and Fusion.
@@
3.2 3D Semantic Network
In addition to projecting 2D pre-trained features onto 3D Gaussians, we alternatively explore a more direct approach – predicting the semantic components from the raw 3D Gaussians. In this section, we build a 3D semantic network $f^{\text{3D}}$ to do exactly that. Specifically, given the input 3D Gaussians $\mathbf{G}$, our 3D network $f^{\text{3D}}$ predict the point-wise semantic component $s^{\text{3D}}$, which can be formulated as Eqn. 1:
@@
We use fused features from Sec. 3.1 (denoted by $s^{\text{2D}}$) to supervise the 3D model. The loss function is the cosine similarity loss:
@@
Though the supervised target entirely comes from pre-trained 2D encoders, the 3D model recognizes the scene by processing 3D geometric information rather than multiple 2D views, making the result more consistent. Experiments in Sec. 4.2.1 show that the 3D prediction $s^{\text{3D}}$ can complement 2D projection features $s^{\text{2D}}$. Also, the inference speed of our 3D semantic model is much faster than 2D projection.
3.3 Inference

### v2

# Semantic Gaussians: Open-Vocabulary Scene Understanding with 3D Gaussian Splatting

@@ Abstract
Open-vocabulary 3D scene understanding presents a significant challenge in computer vision, with wide-ranging applications in embodied agents and augmented reality systems. Existing methods adopt neurel rendering methods as 3D representations and jointly optimize color and semantic features to achieve rendering and scene understanding simultaneously. In this paper, we introduce Semantic Gaussians, a novel open-vocabulary scene understanding approach based on 3D Gaussian Splatting. Our key idea is to distill knowledge from 2D pre-trained models to 3D Gaussians. Unlike existing methods, we design a versatile projection approach that maps various 2D semantic features from pre-trained image encoders into a novel semantic component of 3D Gaussians, which is based on spatial relationship and need no additional training. We further build a 3D semantic network that directly predicts the semantic component from raw 3D Gaussians for fast inference. The quantitative results on ScanNet segmentation and LERF object localization demonstates the superior performance of our method. Additionally, we explore several applications of Semantic Gaussians including object part segmentation, instance segmentation, scene editing, and spatiotemporal segmentation with better qualitative results over 2D and 3D baselines, highlighting its versatility and effectiveness on supporting diverse downstream tasks.
Index Terms:
open-vocabulary scene understanding, 3D Gaussian Splatting.
@@ I Introduction
Various methods have been proposed to achieve open-vocabulary 3D scene understanding, relying on different 3D scene representations such as multi-view RGB images [1], point clouds [2, 3], and Neural Radiance Fields (NeRFs) [4, 5]. Approaches based on these representations have their pros and cons: Multi-view images are the most straightforward representation of 3D scenes, but allowing open-vocabulary understanding usually involves 2D vision-language models [6], which could struggle with consistency across different views, likely due to a lack of visual geometric knowledge; Point clouds are popular and well-studied, but the inherent sparsity nature of point clouds limits the application of open-vocabulary scene understanding upon them  [2], e.g., it is challenging to obtain a dense prediction on a 2D view; injecting open-vocabulary semantics to NeRFs could enjoy both high 2D rendering quality and dense free-form visual recognition [5], but the implicit design requires open-vocabulary recognition training for every new scene, and the rendering speed becomes a bottleneck to achieve real-time high-quality scene understanding.
A recent alternative scene representation is 3D Gaussian Splatting (3DGS) proposed by Kerbl et al. [7]. It utilizes 3D Gaussian points with color, opacity, and covariance matrix to represent the 3D scene, which can be learned from multi-view RGB images via gradient descent training. It attains the NeRF-level view rendering quality while preserving the explicit point-based characteristics similar to point clouds, making it suitable for open-vocabulary 3D scene understanding.
A main branch of previous approaches [8, 9, 5, 10, 4, 11, 12, 13, 14, 15, 16] is to adopt neural rendering methods like NeRF or 3DGS as the 3D representations, and jointly optimizing the color components and the semantic features, to achieve high-quality rendering and 3D scene understanding from arbitrary 2D views. The semantic knowledge is usually distilled from open-vocabulary 2D foundation models, such as CLIP [17] or LSeg [18], whose outputs predicted on training views serve as weak supervision during optimization.
@@
In this work, we propose Semantic Gaussians, a novel approach to open-vocabulary 3D scene understanding building upon the benefits of 3D Gaussian Splatting. The core idea of Semantic Gaussians is to distill the knowledge from pre-trained 2D encoders into 3D Gaussians, thereby assigning a semantic component to each Gaussian point. To achieve this, we establish correspondence between 2D pixels and 3D Gaussian points and propose a versatile projection framework to map the semantic features of 2D pixels onto each 3D Gaussian point. Our framework is rather flexible and can leverage arbitrary pre-trained 2D models, such as OpenSeg [19], CLIP [17], VLPart [20], etc., to generate pixel-wise semantic features on 2D RGB images. Compared to previous approachs, our method injects semantic components into 3D Gaussians without additional training, allowing for effective open-vocabulary scene queries.
@@
In addition to projection, we further introduce a 3D semantic network that directly predicts open-vocabulary semantic components out of raw 3D Gaussians. Specifically, we employ MinkowskiNet [21], a 3D sparse convolution network to process 3D Gaussians. The 3D convolution network takes raw RGB Gaussians as input and is supervised by the semantic components of Gaussians obtained from the aforementioned projection method. As a result, we may simply run this network to obtain the semantic components, enabling faster inference. This network leverages geometric attributes to understand unseen scenes, boosting the generalizability and robustness of our method beyond 2D projection. Note that the prediction of the 3D semantic network can be combined with the projected features to further improve the quality of semantic components in Gaussians and open-vocabulary scene understanding performances.
@@
We conduct experiments on the ScanNet semantic segmentation benchmark [22] and LERF localization [5], and prove our efficiency compared to 2D pre-trained models. Besides segmentation and localization, we also explore diverse applications of Semantic Gaussians, including 3D part segmentation on the MVImgNet object dataset [23], instance segmentation and scene editing in multi-object scenes, and spatiotemporal tracking on 4D dynamic Gaussians [24].
@@
    We conduct experiments on the ScanNet and LERF localization datasets to demonstrate the effectiveness of our method on open-vocabulary scene understanding and explore various applications including object part segmentation, instance segmentation, scene editing, and spatiotemporal tracking.
II Related Work
II-A 3D Scene Representation
@@ II Related Work
Recently, Kerbl et al. [7] have proposed a new novel view synthesis method called 3D Gaussian Splatting, which represents the 3D scene with a set of 3D Gaussians. This method has demonstrated real-time rendering capabilities at 1080p resolution, achieving a remarkable 60 frames per second while maintaining state-of-the-art visual quality. The innovation behind 3D Gaussian Splatting lies in its incorporation of point-based $\alpha$-blending and a differentiable tile rasterizer, enabling efficient rendering. Though it obtains a dense set of Gaussians via optimization, all parameters in these 3D Gaussians are explicit and editable. The speed enhancement and explicit parameterization position 3D Gaussian Splatting as a highly promising representation method. Building upon its success in novel view synthesis, some studies [24, 30, 31, 32, 33, 34] have extended 3DGS to dynamic scenes. For example, Luiten et al. [24] extended the concept to Dynamic 3D Gaussians, explicitly modeling 3D Gaussians at different time steps to accommodate 4D dynamic scenes. Furthermore, many recent works [35, 36, 37, 38, 39, 33, 34] leverage 3D Gaussian Splatting to achieve high-quality text-to-3D or image-to-3D generation. In this study, we propose an open-vocabulary 3D scene understanding method, leveraging the advantages offered by 3D Gaussian Splatting.
II-B Open-Vocabulary Scene Understanding
II-B1 Scene understanding from 2D
@@
Encouraged by the availability of adequate text-image datasets and the advancement in vision language models, the field of 2D open-vocabulary scene understanding has made significant progress in recent years. Prevailing approaches [40, 41, 18] distill knowledge from large-scale pre-trained foundation models (e.g., CLIP [17]) to achieve zero-shot understanding, including recognizing long-tail objects and understanding synonymous labels. However, these methods are limited to small partial scenes represented by a single 2D image. When it comes to 3D scenes, the prediction result of these 2D models can hardly remain consistent between different angles of view. In contrast, our work relies on these 2D pre-trained models to achieve 3D scene understanding, segmenting, and understanding the scene from a panoptic perspective. Moreover, the proposed 3D network can perform 3D-only scene understanding in the absence of 2D pre-trained models and images.
II-B2 Scene understanding from 3D
@@
Open-vocabulary 3D scene understanding has been a long-standing challenge in computer vision. Some point-cloud-based methods[2, 42, 43, 44] encode the semantic features from 2D pre-trained models into 3D scene points to achieve open-vocabulary 3D scene understanding. To achieve high-quality rendering and scene understanding simultaneously, feature field distillation in NeRF has been well explored. Early works such as Semantic-NeRF [26], Panoptic Lifting [45] and Contrastive Lift [46] embed semantic labels into NeRF, resulting in precise 3D segmentation maps. Encouraged by this idea, another branch of methods [8, 9, 5, 10, 4, 11] integrate semantic embeddings from pre-trained models such as LSeg [18], CLIP or DINO [47] into NeRFs, achieving open-vocabulary 3D scene understanding. Recently, some works [12, 13, 14, 15, 16] have made efforts to transfer those NeRF-based methods to 3DGS, obtaining 3DGS with semantic features via optimization. Our work shares a similar idea with those methods, while the Semantic Gaussians requires no extra training for Gaussians. The explicit nature of 3D Gaussian Splatting enables Semantic Gaussians to achieve versatile projection from 2D semantic maps into 3D Gaussian points.
III Semantic Gaussians
@@
In this section, we illustrate the framework of our Semantic Gaussians. Fig. 2 depicts the overall framework of our method. Semantic Gaussians starts from a group of 3D Gaussians (Sec. III-A), performing scene understanding on it through 2D versatile projection and 3D semantic network processing. We first introduce our versatile projection method that projects 2D semantic embeddings from various pre-trained vision-language models into 3D Gaussian points (Sec. III-B). We then depict the 3D semantic network that learns from the projected features and predicts the semantics of 3D Gaussians in unseen scenes (Sec. III-C). At last, we describe the feature ensemble process and use the prediction result to support various applications (Sec. III-D).
III-A 3D Gaussian Splatting
To achieve general 3D open-vocabulary scene understanding, we employ 3D Gaussian Splatting [7] as the representation of 3D scenes. 3DGS can render images from arbitrary viewpoints in a differentiable manner, thus effectively leveraging the knowledge of various 2D foundation models. Specifically, we achieve scene understanding by rendering 2D semantic images from specified viewpoints with 3DGS.
3DGS consists of a set of learnable 3D Gaussian points, where each point has a 3D coordinate $\mu$ representing its position, a covariance matrix $\Sigma$ representing its shape, spherical harmonic parameters $c$ representing its color, and an opacity value $\alpha$ representing its transparency. 3DGS can be constructed from multi-view images and can utilize information from Structure-from-Motion (SfM) point clouds [48] for initialization, thereby achieving better rendering quality and geometric structure.
3DGS uses point-based $\alpha$-blending to compute pixel values on 2D images. The value of each pixel $C$ is given by volumetric rendering along a ray:
$C=\sum_{i\in\mathcal{N}}c_{i}\alpha_{i}T_{i}\text{ with }T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (1)
where $\mathcal{N}$ is the set of sorted Gaussians in front-to-back depth order overlapping with the given pixel. $\alpha_{i}$ is the opacity of point $i$, and $c_{i}$ is the color of point $i$, which is calculated by spherical harmonics.
To project 3D Gaussians onto a certain 2D plane, Zwicker et al. proposed a splatting method to calculate the covariance matrix $\Sigma^{\prime}$ from the camera’s viewpoint. Given the world-to-camera transformation matrix $W$, the covariance matrix in camera coordinates is given as follows:
$\Sigma^{\prime}=JW\Sigma W^{T}J^{T},$ (2)
where J is the Jacobian of the affine approximation of the projective transformation. If we skip the third row and column of $\Sigma^{\prime}$, we can obtain the 2D variance matrix. To assure the positive semi-definite, the covariance matrix $\Sigma$ is decomposed into rotation matrix $R$ and scaling matrix $S$, which can be represented as follows:
$\Sigma=RSS^{T}R^{T}.$ (3)
3DGS can be regarded as a special point cloud with additional features, thus sharing some properties of point clouds. An intuitive idea is to project the semantic information obtained from 2D foundation models onto the corresponding Gaussian points based on spatial relationships, rather than through differentiable rasterization and rendering. When rendering semantic maps, it is sufficient to consider the correspondence of geometric positions without accounting for complex lighting conditions. Based on this idea, we propose Semantic Gaussians to achieve versatile scene understanding.
III-B 2D Versatile Projection
@@ III-B1 Semantic Map Extraction
III-B2 Unifiying Various 2D Features with SAM
Accommodating a variety of 2D pre-trained features is non-trivial as they can be pixel-level segmentation network (e.g., OpenSeg [19], LSeg [18]), instance-level recognition network (e.g., GroundingDINO [49], VLPart [20]), or image-level classification network (e.g., CLIP [17]). Additionally, Semantic Gaussians are able to utilizes Segment Anything (SAM) [50] to produce fine segmentation maps for each model. For pixel-level models, SAM can refine the segmentation boundary. Given an RGB image $\mathbf{I}$, we use everything prompt in SAM to generate $N$ binary masks $\mathbf{M}_{1},\cdots,\mathbf{M}_{N}$. We calculate the average pooling of embeddings in each mask $\mathbf{M}_{i}$, and assign it as the embedding of all pixels in this mask: $s[\mathbf{M}_{i}]=AvgPool(s[\mathbf{M}_{i}])$. For instance-level models, we use SAM as the postprocessing module to get fine masks. Similar to Grounded-SAM [51], we use the prediction result from the pre-trained model as the box prompt of SAM. After getting the binary mask, we assign the CLIP embedding of this instance to all the pixels within the instance region. For image-level models, SAM can be a preprocessing module to get region proposals. We use the “everything” prompt in SAM to get various proposed regions. Each region is padded, cropped, and resized to $224\times 224$, and fed into the image-level model to get semantic embeddings. Similarly, the semantic embeddings are assigned to all pixels within each proposed region.
III-B3 2D-3D Projection and Fusion
@@ III-B2 Unifiying Various 2D Features with SAM
Figure 3: Visualization of scene-level semantic segmentation performance for open-vocabulary 3D scene understanding methods on ScanNet dataset.
III-C 3D Semantic Network
In addition to projecting 2D pre-trained features onto 3D Gaussians, we alternatively explore a more direct approach – predicting the semantic components from the raw 3D Gaussians. In this section, we build a 3D semantic network $f^{\text{3D}}$ to do exactly that. Specifically, given the input 3D Gaussians $\mathbf{G}$, our 3D network $f^{\text{3D}}$ predict the point-wise semantic component $s^{\text{3D}}$, which can be formulated as Eqn. 4:
@@
We use fused features from Sec. III-B (denoted by $s^{\text{2D}}$) to supervise the 3D model. The loss function is the cosine similarity loss:
@@
Though the supervised target entirely comes from pre-trained 2D encoders, the 3D model recognizes the scene by processing 3D geometric information rather than multiple 2D views, making the result more consistent. Experiments in Sec. IV-B1 show that the 3D prediction $s^{\text{3D}}$ can complement 2D projection features $s^{\text{2D}}$. Also, the inference speed of our 3D semantic model is much faster than 2D projection.
III-D Inference